OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

You want to use the formula: y-y1=m(x-x1) and plug in your x and y values

OpenStudy (sidsiddhartha):

\[y=mx+c\] y=-4 ,x=2 and m=3/5 so calculate c and then substitute it in the main equation

OpenStudy (anonymous):

2.6?

OpenStudy (sidsiddhartha):

-4=(3/5)*2+c c=-26/5=-5.2

OpenStudy (haseeb96):